How do you work out a problem in the form of 3 log base 5 n-2 log base 5 5n = log5 7?
Answer by lwsshak3(11628)

How do you work out a problem in the form of 3 log base 5 n-2 log base 5 5n = log5 7?
**
3log5(n)-2log5(5n)=log5(7)
3log5(n)-2log5(5n)-log5(7)=0
3log5(n)-2(log5(5)+log5(n))-log5(7)=0
3log5(n)-2(1+log5(n))-log5(7)=0
3log5(n)-2-2log5(n)-log5(7)=0
log5(n)-log5(7)=2
log5[n/7]=2
convert to exponential form
5^2=n/7=25
n=7*25=175